OpenStudy (samigupta8):

A ball of mass m falls vertically from a height h and collides with a block of equal mass m moving horizontally with a velocity v on a surface.The coefficient of kinetic friction between thr block and the surface is 0.2,while the coefficient of restitution between the ball and block is 0.5.There is no friction acting between the block and the ball.The velocity of block just after collision decreases by

OpenStudy (samigupta8):

@michele_laino

OpenStudy (samigupta8):

@ganeshie8

OpenStudy (samigupta8):

I m not getting how can this be possible m√2gh=mVy And from thr coefficient of restitution i got this e=-Vy/Uy..

OpenStudy (samigupta8):

I got this first equation by momentum conservation along y axis..

ganeshie8 (ganeshie8):

.

OpenStudy (michele_laino):

I think to have the solution, please tag, so we can start the corresponding discussion

OpenStudy (pawanyadav):

You can get another equation by conservation of linear momentum along x-axis.

OpenStudy (samigupta8):

@michele_laino

OpenStudy (samigupta8):

And how can we apply momentum conservation along x axis when we know that there is a force in that direction for the sytem concerned.. (Friction force)

OpenStudy (michele_laino):

there is an impulsive force between the two objects which takes place during the collision

OpenStudy (michele_laino):

such force \(F\) does exist for a short time interval \(\Delta t\)

OpenStudy (samigupta8):

Do u wanna say that mu-integral of Ndt =mvy

OpenStudy (samigupta8):

Normla force is providing the necessary impluse to the ball here..

OpenStudy (michele_laino):

no, please note that now, the transferred momentum, is: \[\Delta Q = \frac{{m{v_0}}}{2} = m\sqrt {\frac{{gh}}{2}} \]

OpenStudy (michele_laino):

since the coefficient of restitution is \(1/2\)

OpenStudy (michele_laino):

so, I apply the theorem of conservation of kinetic energy, and I can write: \[\frac{{m{V^2}}}{2} - \mu \left( {mg + F} \right)\Delta tV = \frac{{mV_1^2}}{2}\]

OpenStudy (michele_laino):

where \(V,\;V_1\) are the velocities of the horizontal moving block before, and after the collision respectively

OpenStudy (samigupta8):

F??

OpenStudy (michele_laino):

\(F\) is the impulsive force, which take place during collision, between the two blocks

OpenStudy (michele_laino):

takes*

OpenStudy (michele_laino):

next, we can make these approximations: \[\begin{gathered} F \gg mg \hfill \\ \hfill \\ V \cong {V_1} \hfill \\ \end{gathered} \]

OpenStudy (samigupta8):

Sir do u mean to say that it is due to impulse force F that the ball after collision will move in x direction also..??

OpenStudy (michele_laino):

more precisely \(F\) takes account of the weight of the vertical moving body

OpenStudy (michele_laino):

yes! Since the collision prevents the horizontally moving body to change is vertical coordinate, which is, and it will be zero

OpenStudy (michele_laino):

its*

OpenStudy (samigupta8):

If it takes into account thr vertical motion then how can we apply this . 1/2mv^2- u(mg+F)∆t=1/2mv'^2 U wrote that impulsive force is in vertical direction n u r applying it for horizontal motion of block .. How??

OpenStudy (michele_laino):

since the impulsive force causes an increasing of the friction force, which is horizonatlly directed

OpenStudy (samigupta8):

Sir it should be Mv-integral of u(mg-F)dt=Mv'

OpenStudy (samigupta8):

Sorry mg+F

OpenStudy (michele_laino):

If I neglect \(mg\), then \(F\;\Delta t\) can be considered, with a good approximation, the momentum transferred during collision now we have the subsequent steps: \[\begin{gathered} \frac{m}{2}\left( {{V^2} - V_1^2} \right) = \mu \left( {mg + F} \right)\Delta tV \hfill \\ \hfill \\ \frac{m}{2}\left( {V + {V_1}} \right)\left( {V - {V_1}} \right) = \mu \left( {mg + F} \right)\Delta tV \hfill \\ \hfill \\ \frac{m}{2} \cdot 2V\left( {V - {V_1}} \right) = \mu F\Delta tV \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

therefore, I got this velocity change: \[V - {V_1} = \frac{{\mu F\Delta t}}{m} = \mu \sqrt {\frac{{gh}}{2}} \]

OpenStudy (samigupta8):

I got v-v1 as mu√gh/2

OpenStudy (michele_laino):

it is right then!

OpenStudy (samigupta8):

That gives us the ans as 0.1√2gh

OpenStudy (samigupta8):

Correct??

OpenStudy (michele_laino):

yes! If we rationalize, we get: \[V - {V_1} = \frac{{\mu F\Delta t}}{m} = \mu \sqrt {\frac{{gh}}{2}} = \frac{{0.2\sqrt 2 }}{2}\sqrt {gh} = 0.1\sqrt 2 \sqrt {gh} \]

OpenStudy (samigupta8):

Bt ans is diff in ans book..

OpenStudy (samigupta8):

They say it's 0.3√2gh

OpenStudy (samigupta8):

Gudnyt sir!! Gotto go now!

OpenStudy (michele_laino):

ok! goodnight!

OpenStudy (samigupta8):

If you get the ans they gave then pls do inform me